# Understanding LLM Development Through Longitudinal Study: Insights from the Open Ko-LLM Leaderboard

### v1

@@ Abstract
Chanjun Park, Hyeonwoo Kim Upstage AI {chanjun.park, choco_9966}@upstage.ai
@@ 2.3 Temporal Shifts in Leaderboard Ranking Patterns
Task Correlations Over Time.
Performance Trends by Model Type.
Performance Trends Across Model Sizes.
3 Conclusion
This study provides a longitudinal analysis of the Open Ko-LLM Leaderboard, revealing key performance trends: smaller models face scalability limitations, while larger models experience saturation without advancements in pretrained models. These findings highlight the need for continuous innovation in pretrained models to enhance LLM capabilities. Additionally, the study shows that analyzing leaderboard data offers valuable insights into the evolving dynamics of LLM performance.
Acknowledgments
Limitations
Ethics Statement
Appendix A Open Ko-LLM Leaderboard
Appendix B Additional Analysis
@@
Figure 6 presents the monthly distribution of submissions across different model types on the Open Ko-LLM leaderboard. Initially, pretrained models constituted 37% of all submissions, but this proportion declined sharply over time, with no pretrained models submitted by August 2024. This trend signals a diminishing focus on pretrained models within the community, which is concerning given their foundational importance discussed in Section 2. Therefore, a renewed emphasis on fostering interest and engagement with pretrained models could help address this emerging gap.

### v2

@@ Abstract
1Chanjun Park, 2Hyeonwoo Kim 1Korea University, 2Upstage AI bcj1210@korea.ac.kr choco_9966@upstage.ai
@@ 3.3 Temporal Shifts in Leaderboard Ranking Patterns
Task Correlations Over Time.
Performance Trends by Model Type.
Performance Trends Across Model Sizes.
3.4 Evaluation Patterns and Submission Insights
Figure 6 presents the monthly distribution of submissions across different model types on the Open Ko-LLM leaderboard. Initially, pretrained models constituted 37% of all submissions, but this proportion declined sharply over time, with no pretrained models submitted by August 2024. This trend signals a diminishing focus on pretrained models within the community, which is concerning given their foundational importance discussed in Section 3.3. Therefore, a renewed emphasis on fostering interest and engagement with pretrained models could help address this emerging gap.
4 Conclusion
This study provides a longitudinal analysis of the Open Ko-LLM Leaderboard, uncovering significant performance trends and underlying challenges in LLM development. It was observed that smaller models consistently face scalability limitations, preventing substantial performance advancements. In contrast, larger models initially show promising improvements but eventually reach a saturation point, highlighting a critical dependency on advancements in pretrained models. These findings underscore the need for continuous innovation and enhancement in the development of pretrained models to push the boundaries of LLM capabilities further. Additionally, the analysis demonstrates the utility of leaderboard data in tracking the evolving dynamics of LLM performance. By examining a broader range of model submissions and evaluation patterns over an extended period, this study provides valuable insights into how model size, type, and tuning methods influence overall effectiveness. Such insights can inform targeted research efforts and encourage the development of strategies aimed at overcoming existing limitations, ultimately supporting more robust and adaptable LLMs.
Acknowledgments
Limitations
Ethics Statement